OpenStudy (anonymous):

A steel rail is built using 25 m tracks at (25 degrees celsius). How wide should the expansion slots between the tracks at 25 degrees celsius to prevent buckling if the range of temperature is 18 to 40 degrees celsius? Answer in a separate sheet.

OpenStudy (john_es):

Do you know the linear thermal expansion coefficient of the steel?

OpenStudy (john_es):

Well, suppose the expansion coefficient is alpha, then, you should calculate the ranges, \[\Delta L=\alpha L(T_1-T_0)\]\[\Delta L=\alpha L(T_2-T_0)\] Supposing, T_0=25 degrees. This will give you the answer.

OpenStudy (anonymous):

should L = 25 m?, so the two ΔL = -2.1 x 10 ^(-3) and 4.5 x 10 ^(-3)?

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

Yes, the lengths are of these order. I don't know the exact coefficient that you have used in your problem, but I would suppose is the correct one.

OpenStudy (john_es):

Remember that you only need the hottest temp at the end. That is your answer should be the maximum (positive) length.

OpenStudy (anonymous):

@John_ES , thank you so much :)